OpenStudy (anonymous):

find the product of : x(3x-7)(x+4)

OpenStudy (aroub):

jut distribute :)

OpenStudy (anonymous):

Start with the two binomial terms and use FOIL (First, Outer, Inner, Last)\[(3x-7)(x+4)=3x^{2}+12x-7x-28=3x^{2}+5x-28\] So you now have \[x(3x^{2}+5x-28)\] Using distribution, you get \[x(3x^{2}+5x-28)=3x^{3}+5x^{2}-28x\]

OpenStudy (anonymous):

x(3x-7)(x+4) = (3x*x) + (3x*4)+(-7*x)+(-7*4)=(3x^2 + 12x - 7x - 28)*x = => 3x^3+12x^2-7x^2-28x